# Rigidity results for initial data sets satisfying the dominant energy condition

### v3

@@ 4.4. Proof of Theorem I in the simply connected case
We define $\mathcal{W}=\mathcal{S}\oplus\mathcal{S}$. We define a connection $\nabla^{\mathcal{W}}$ on $\mathcal{W}$ by
@@
For each point $p\in M$, we define a map $\gamma^{\mathcal{W}}:\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$ by
@@
where $X\in TM$ and $u,v\in\mathcal{S}$.
@@
The map $\gamma^{\mathcal{W}}:\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$ satisfies the Clifford relations with respect to the Lorentzian inner product on $\mathbb{R}^{n,1}$. In other words,
$\gamma^{\mathcal{W}}(a)\,\gamma^{\mathcal{W}}(b)+\gamma^{\mathcal{W}}(b)\,\gamma^{\mathcal{W}}(a)=-2\,\langle a,b\rangle_{\mathbb{R}^{n,1}}\,\text{\rm id}$
@@
Note that
$\gamma^{\mathcal{W}}(X)\,\gamma^{\mathcal{W}}(Y)+\gamma^{\mathcal{W}}(Y)\,\gamma^{\mathcal{W}}(X)=-2\,g(X,Y)\,\text{\rm id}$
@@
for all $X\in TM$ and
$(\gamma^{\mathcal{W}}(T))^{2}=\text{\rm id}.$
This completes the proof of Lemma 4.22. ∎
@@
We decompose $a=Y+fT$, where $Y$ is a section of $TM$ and $f$ is a scalar function on $M$. Then $\bar{D}_{X}Y+\bar{D}_{X}(fT)=0$, where $\bar{D}$ denotes the standard flat connection on $\mathbb{R}^{n,1}$. Using the formula $\bar{D}_{X}Y=D_{X}Y+q(X,Y)\,T$, we obtain
@@
$-\gamma^{\mathcal{W}}(D_{X}Y)\,(u,v)-q(X,Y)\,\gamma^{\mathcal{W}}(T)\,(u,v)$
@@
$-(\gamma(D_{X}Y)\,v,\gamma(D_{X}Y)\,u)-q(X,Y)\,(-iv,iu)$
$=0$
for all $X,Y\in TM$. Using the formula $\bar{D}_{X}T=q(X)$, we obtain
@@
$=0$
for all $X\in TM$. This implies
(4.7) $\nabla_{X}^{\mathcal{W}}(\gamma^{\mathcal{W}}(fT)\,(u,v))-\gamma^{\mathcal{W}}(fT)\,\nabla_{X}^{\mathcal{W}}(u,v)-\gamma^{\mathcal{W}}(\bar{D}_{X}(fT))\,(u,v)=0$
for all $X\in TM$. In the next step, we add (4.4) and (4.7). Using the identity $\bar{D}_{X}Y+\bar{D}_{X}(fT)=0$, we conclude that
$\nabla_{X}^{\mathcal{W}}(\gamma^{\mathcal{W}}(a)\,(u,v))-\gamma^{\mathcal{W}}(a)\,\nabla_{X}^{\mathcal{W}}(u,v)=0$
for all $X\in TM$. This completes the proof of Lemma 4.23. ∎
@@
For each point $p\in M$, we define a map $\gamma^{\mathcal{W}}:\Lambda^{2}\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$ by
$\gamma^{\mathcal{W}}(a\wedge b)=\frac{1}{2}\,(\gamma^{\mathcal{W}}(a)\,\gamma^{\mathcal{W}}(b)-\gamma^{\mathcal{W}}(b)\,\gamma^{\mathcal{W}}(a)).$
@@
Suppose that $p$ is a point in $M$ and $\omega\in\Lambda^{2}\mathbb{R}^{n,1}$ is a two-form with the property that $\gamma^{\mathcal{W}}(\omega)=0\in\text{\rm End}(\mathcal{W}|_{p})$. Then $\omega=0$.
@@
We decompose $\omega=\mu+T\wedge\eta$, where $\mu\in\Lambda^{2}T_{p}M$ and $\eta\in T_{p}M$. By assumption,
00 0 $=\gamma^{\mathcal{W}}(\omega)\,(u,v)$
$=\gamma^{\mathcal{W}}(\mu)\,(u,v)+\gamma^{\mathcal{W}}(T)\,\gamma^{\mathcal{W}}(\eta)\,(u,v)$
$=(\gamma(\mu)\,u-i\,\gamma(\eta)\,u,\gamma(\mu)\,v+i\,\gamma(\eta)\,v)$
for all $u,v\in\mathcal{S}|_{p}$. Therefore, $\gamma(\mu)-i\,\gamma(\eta)=0\in\text{\rm End}(\mathcal{S}|_{p})$ and $\gamma(\mu)+i\,\gamma(\eta)=0\in\text{\rm End}(\mathcal{S}|_{p})$. This implies $\gamma(\mu)=0\in\text{\rm End}(\mathcal{S}|_{p})$ and $\gamma(\eta)=0\in\text{\rm End}(\mathcal{S}|_{p})$. Thus, $\mu=0$ and $\eta=0$. This completes the proof of Lemma 4.25. ∎
By Proposition 4.1, there exists a $\nabla^{\mathcal{E},q}$-parallel section $s$ of the bundle $\text{\rm Hom}(S_{0},\mathcal{S})$ with the property that $s$ is invertible at each point in $M$. Moreover, $s$ satisfies the boundary condition $\gamma(\nu)\circ s=s\circ\gamma_{0}(N)$ at each point on the boundary $\Sigma$.
Replacing $q$ by $-q$, we can find a $\nabla^{\mathcal{E},-q}$-parallel section $t$ of the bundle $\text{\rm Hom}(S_{0},\mathcal{S})$ with the property that $t$ is invertible at each point in $M$. Moreover, $t$ satisfies the boundary condition $\gamma(\nu)\circ t=t\circ\gamma_{0}(N)$ at each point on the boundary $\Sigma$.
We denote by $\mathcal{Y}$ the set of all pairs $(v,\omega)\in\mathbb{R}^{n}\times\Lambda^{2}\mathbb{R}^{n,1}$ with the property that
$i\,\gamma^{\mathcal{W}}(\omega)\circ(s,t)=(s\circ\gamma_{0}(v),-t\circ\gamma_{0}(v))$
at each point on $M$. Clearly, $\mathcal{Y}$ is a linear subspace of $\mathbb{R}^{n}\times\Lambda^{2}\mathbb{R}^{n,1}$. Let $\mathcal{Z}\subset\Lambda^{2}\mathbb{R}^{n,1}$ denote the image of $\mathcal{Y}$ under the projection $\mathcal{Y}\to\Lambda^{2}\mathbb{R}^{n,1},\,(v,\omega)\mapsto\omega$.
@@
We have $\mathcal{Y}=\{\mathbb{R}\cdot(N(p),T(p)\wedge\nu(p)):p\in\Sigma\}$. The projection $\mathcal{Y}\to\mathbb{R}^{n},\,(v,\omega)\mapsto v$ is bijective. In particular, $\mathcal{Y}$ has dimension $n$.
@@
Using the boundary conditions satisfied by $s$ and $t$, we compute
$i\,\gamma^{\mathcal{W}}(T\wedge\nu)\circ(s,t)$ $=i\,\gamma^{\mathcal{W}}(T)\,\gamma^{\mathcal{W}}(\nu)\circ(s,t)$
$=(\gamma(\nu)\circ s,-\gamma(\nu)\circ t)$
$=(s\circ\gamma_{0}(N),-t\circ\gamma_{0}(N))$
at each point on $\Sigma$. Let us fix a boundary point $p\in\Sigma$, and let $v=N(p)\in\mathbb{R}^{n}$, $a=T(p)\in\mathbb{R}^{n,1}$, and $b=\nu(p)\in\mathbb{R}^{n,1}$. Then
$i\,\gamma^{\mathcal{W}}(a\wedge b)\circ(s,t)=(s\circ\gamma_{0}(v),-t\circ\gamma_{0}(v))$
at the point $p$. Since the left hand side and the right hand side are parallel sections with respect to the same connection, we conclude that
$i\,\gamma^{\mathcal{W}}(a\wedge b)\circ(s,t)=(s\circ\gamma_{0}(v),-t\circ\gamma_{0}(v))$
at each point in $M$. Thus, $(v,a\wedge b)\in\mathcal{Y}$. To summarize, we have shown that
$\{\mathbb{R}\cdot(N(p),T(p)\wedge\nu(p)):p\in\Sigma\}\subset\mathcal{Y}.$
Since the map $N:\Sigma\to S^{n-1}$ is surjective, it follows that the projection $\mathcal{Y}\to\mathbb{R}^{n},\,(v,\omega)\mapsto v$ is surjective. On the other hand, it follows from Lemma 4.25 that the projection $\mathcal{Y}\to\mathbb{R}^{n},\,(v,\omega)\mapsto v$ is injective. Putting these facts together, we conclude that
$\{\mathbb{R}\cdot(N(p),T(p)\wedge\nu(p)):p\in\Sigma\}=\mathcal{Y}.$
This completes the proof of Proposition 4.26. ∎
@@
We have $\mathcal{Z}=\{\mathbb{R}\cdot(T(p)\wedge\nu(p)):p\in\Sigma\}$. Moreover, $\mathcal{Z}$ has dimension $n$.
@@
The first statement follows directly from Proposition 4.26. The second statement follows from the fact that $\mathcal{Y}$ has dimension $n$ and the projection $\mathcal{Y}\to\Lambda^{2}\mathbb{R}^{n,1},\,(v,\omega)\mapsto\omega$ is injective. ∎
@@
By Theorem 6 in [LA], there are two possibilities:
Case 1: We have $n=3$ and we can find three linearly independent vectors $b_{0},b_{1},b_{2}\in\mathbb{R}^{3,1}$ with the property that $\mathcal{Z}=\text{\rm span}\{b_{0}\wedge b_{1},b_{1}\wedge b_{2},b_{2}\wedge b_{0}\}$. We can find two linearly independent vectors $\tilde{b}_{1},\tilde{b}_{2}\in\text{\rm span}\{b_{0},b_{1},b_{2}\}$ such that $\text{\rm span}\{\tilde{b}_{1},\tilde{b}_{2}\}$ is spacelike. Then $\tilde{b}_{1}\wedge\tilde{b}_{2}\in\mathcal{Z}$. By Corollary 4.27, we can find a point $p\in\Sigma$ with the property that $\tilde{b}_{1}\wedge\tilde{b}_{2}$ is a non-zero scalar multiple of $T(p)\wedge\nu(p)$. In particular, $\text{\rm span}\{\tilde{b}_{1},\tilde{b}_{2}\}=\text{\rm span}\{T(p),\nu(p)\}$. This contradicts the fact that $\text{\rm span}\{\tilde{b}_{1},\tilde{b}_{2}\}$ does not contain a timelike vector.
Case 2: There exists a basis $\{b_{0},b_{1},\ldots,b_{n}\}$ of $\mathbb{R}^{n,1}$ with the property that $\mathcal{Z}=\text{\rm span}\{b_{0}\wedge b_{1},b_{0}\wedge b_{2},\ldots,b_{0}\wedge b_{n}\}$. We claim that $b_{0}$ is timelike. Suppose that $b_{0}$ is spacelike or null. We can find a vector $\tilde{b}\in\mathbb{R}^{n,1}$ such that $\tilde{b}$ is orthogonal to $b_{0}$ and $\tilde{b}$ is spacelike. This implies that $b_{0}$ and $\tilde{b}$ are linearly independent. Moreover, $b_{0}\wedge\tilde{b}\in\mathcal{Z}$. By Corollary 4.27, we can find a point $p\in\Sigma$ with the property that $b_{0}\wedge\tilde{b}$ is a non-zero scalar multiple of $T(p)\wedge\nu(p)$. In particular, $\text{\rm span}\{b_{0},\tilde{b}\}=\text{\rm span}\{T(p),\nu(p)\}$. This contradicts the fact that $\text{\rm span}\{b_{0},\tilde{b}\}$ does not contain a timelike vector. ∎
@@
For each point $p\in\Sigma$, we know that $T(p)\wedge\nu(p)\in\text{\rm span}\{b_{0}\wedge b_{1},b_{0}\wedge b_{2},\ldots,b_{0}\wedge b_{n}\}$, hence $b_{0}\in\text{\rm span}\{T(p),\nu(p)\}$. Consequently, $\langle b_{0},X\rangle_{\mathbb{R}^{n,1}}=0$ for every tangent vector $X\in T\Sigma$. In particular, each connected component of $\Sigma$ is contained in a hyperplane in $\mathbb{R}^{n,1}$ which is orthogonal to $b_{0}$.
For each point $p\in\Sigma$, we can find a spacelike unit vector $B(p)\in\mathbb{R}^{n,1}$ such that $B(p)$ is orthogonal to $b_{0}$ and $T(p)\wedge\nu(p)=b_{0}\wedge B(p)$. This defines a smooth map $B:\Sigma\to S^{n-1}\subset\mathbb{R}^{n,1}$. The map $B$ can be interpreted as the unit normal to $\Sigma$ within the hyperplane orthogonal to $b_{0}$.
Using Proposition 4.26, we obtain $\mathcal{Y}=\{\mathbb{R}\cdot(N(p),b_{0}\wedge B(p)):p\in\Sigma\}$. This implies that the map $B:\Sigma\to S^{n-1}\subset\mathbb{R}^{n,1}$ is the composition of the map $N:\Sigma\to S^{n-1}$ with a linear isometry of $\mathbb{R}^{n}$. In particular, $\|dB\|_{\text{\rm tr}}=\|dN\|_{\text{\rm tr}}$ at each point on $\Sigma$.
@@
If $p$ is a point on the boundary $\Sigma$ with $\langle b_{0},\nu\rangle_{\mathbb{R}^{n,1}}\neq 0$, then $h=0$, $dN=0$, and $dB=0$ at the point $p$.
@@
Let us fix a point $p\in\Sigma$, and let $\{e_{1},\ldots,e_{n-1}\}$ denote an orthonormal basis of $T_{p}\Sigma$. Differentiating the identity $T\wedge\nu=b_{0}\wedge B$ gives

### v4

@@ 4.4. Proof of Theorem I in the simply connected case
We define a vector bundle $\mathcal{W}$ over $M$ by $\mathcal{W}=\mathcal{S}\oplus\mathcal{S}$. We define a connection $\nabla^{\mathcal{W}}$ on $\mathcal{W}$ by
@@
For each point $p\in M$, we define a map $\gamma^{\mathcal{W}}\colon\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$ by
@@
where $X$ denotes a vector field on $M$ and $u,v\in\mathcal{S}|_{p}$.
@@
The map $\gamma^{\mathcal{W}}\colon\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$ satisfies the Clifford relations with respect to the Lorentzian inner product on $\mathbb{R}^{n,1}$. In other words,
$\gamma^{\mathcal{W}}(a)\,\gamma^{\mathcal{W}}(b)+\gamma^{\mathcal{W}}(b)\,\gamma^{\mathcal{W}}(a)=-2\,\langle a,b\rangle_{\mathbb{R}^{n,1}}\,\textup{id}$
@@
We compute
$\gamma^{\mathcal{W}}(X)\,\gamma^{\mathcal{W}}(Y)+\gamma^{\mathcal{W}}(Y)\,\gamma^{\mathcal{W}}(X)=-2\,g(X,Y)\,\textup{id}$
@@
for all $X\in TM$. Finally, $(\gamma^{\mathcal{W}}(T))^{2}=\textup{id}$. This completes the proof of Lemma 4.21. ∎
@@
Let $X$ be a section of $TM$. We decompose $a=Y+fT$, where $Y$ is a section of $TM$ and $f$ is a scalar function on $M$. Then $\bar{D}_{X}Y+\bar{D}_{X}(fT)=0$, where $\bar{D}$ denotes the standard flat connection on $\mathbb{R}^{n,1}$. Using the identity $\bar{D}_{X}Y=D_{X}Y+q(X,Y)\,T$, we obtain
@@
$\quad-\gamma^{\mathcal{W}}(D_{X}Y)\,(u,v)-q(X,Y)\,\gamma^{\mathcal{W}}(T)\,(u,v)$
@@
$\quad-(\gamma(D_{X}Y)\,v,\gamma(D_{X}Y)\,u)-q(X,Y)\,(-iv,iu)$
$=0.$
Using the identity $\bar{D}_{X}T=q(X)$, we obtain
@@
$=0.$
This implies
(4.9) $\nabla_{X}^{\mathcal{W}}(\gamma^{\mathcal{W}}(fT)\,(u,v))-\gamma^{\mathcal{W}}(fT)\,\nabla_{X}^{\mathcal{W}}(u,v)-\gamma^{\mathcal{W}}(\bar{D}_{X}(fT))\,(u,v)=0.$
In the next step, we add (4.4) and (4.9). Using the identity $\bar{D}_{X}Y+\bar{D}_{X}(fT)=0$, we conclude that
$\nabla_{X}^{\mathcal{W}}(\gamma^{\mathcal{W}}(a)\,(u,v))-\gamma^{\mathcal{W}}(a)\,\nabla_{X}^{\mathcal{W}}(u,v)=0.$
This completes the proof of Lemma 4.22. ∎
@@
For each point $p\in M$, we define a map $\gamma^{\mathcal{W}}\colon\Lambda^{2}\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$ by
$\gamma^{\mathcal{W}}(a\wedge b)=\tfrac{1}{2}\,(\gamma^{\mathcal{W}}(a)\,\gamma^{\mathcal{W}}(b)-\gamma^{\mathcal{W}}(b)\,\gamma^{\mathcal{W}}(a)).$
@@
By Proposition 4.1, there exists a $\nabla^{\mathcal{E},q}$-parallel section $s$ of the bundle $\text{\rm Hom}(S_{0},\mathcal{S})$ with the property that $s$ is invertible at each point in $M$. Moreover, $s$ satisfies the boundary condition $\gamma(\nu)\circ s=s\circ\gamma_{0}(N)$ at each point on $\Sigma$.
Replacing $q$ by q-q, we can find a $\nabla^{\mathcal{E},-q}$-parallel section $t$ of the bundle $\text{\rm Hom}(S_{0},\mathcal{S})$ with the property that $t$ is invertible at each point in $M$. Moreover, $t$ satisfies the boundary condition $\gamma(\nu)\circ t=t\circ\gamma_{0}(N)$ at each point on $\Sigma$.
In the following, we fix sections $s$ and $t$ of $\text{\rm Hom}(S_{0},\mathcal{S})$ satisfying the properties above. For each $\sigma\in S_{0}$ and each $\tau\in S_{0}$, the pair $(s\sigma,t\tau)$ defines a $\nabla^{\mathcal{W}}$-parallel section of the vector bundle $\mathcal{W}$.
Definition 4.24.
We denote by $\mathcal{Y}$ the set of all pairs $(\xi,\omega)\in\mathbb{R}^{n}\times\Lambda^{2}\mathbb{R}^{n,1}$ with the property that, for all $\sigma,\tau\in S_{0}$,
$(s\gamma_{0}(\xi)\sigma,-t\gamma_{0}(\xi)\tau)=i\,\gamma^{\mathcal{W}}(\omega)\,(s\sigma,t\tau)$
at each point on $M$. Let $\mathcal{Z}\subset\Lambda^{2}\mathbb{R}^{n,1}$ denote the image of $\mathcal{Y}$ under the projection $\mathcal{Y}\to\Lambda^{2}\mathbb{R}^{n,1},\,(\xi,\omega)\mapsto\omega$.
Clearly, $\mathcal{Y}$ is a linear subspace of $\mathbb{R}^{n}\times\Lambda^{2}\mathbb{R}^{n,1}$, and $\mathcal{Z}$ is a linear subspace of $\Lambda^{2}\mathbb{R}^{n,1}$.
@@
Suppose that $\xi\in\mathbb{R}^{n}$ satisfies $(\xi,0)\in\mathcal{Y}$. Then $\xi=0$.
@@
Let us fix an arbitrary point $p\in M$. Since $(\xi,0)\in\mathcal{Y}$, we obtain
$(s\gamma_{0}(\xi)\sigma,-t\gamma_{0}(\xi)\tau)=(0,0)$
for all $\sigma,\tau\in S_{0}$. Since $s,t\in\text{\rm Hom}(S_{0},\mathcal{S}|_{p})$ are invertible, it follows that $\gamma_{0}(\xi)=0\in\text{\rm End}(S_{0})$. Thus, $\xi=0$. ∎
Lemma 4.26.
Suppose that $\omega\in\Lambda^{2}\mathbb{R}^{n,1}$ satisfies $(0,\omega)\in\mathcal{Y}$. Then $\omega=0$.
Proof.
Fix an arbitrary point $p\in M$. We may write $\omega=\rho+T\wedge\eta$, where $\rho\in\Lambda^{2}T_{p}M$ and $\eta\in T_{p}M$. Since $\langle T,\eta\rangle_{\mathbb{R}^{n,1}}=0$, we obtain $\gamma^{\mathcal{W}}(T\wedge\eta)=\gamma^{\mathcal{W}}(T)\,\gamma^{\mathcal{W}}(\eta)$. Since $(0,\omega)\in\mathcal{Y}$, we compute
$(0,0)$ $=\gamma^{\mathcal{W}}(\omega)\,(s\sigma,t\tau)$
$=\gamma^{\mathcal{W}}(\rho)\,(s\sigma,t\tau)+\gamma^{\mathcal{W}}(T)\,\gamma^{\mathcal{W}}(\eta)\,(s\sigma,t\tau)$
$=(\gamma(\rho)s\sigma-i\,\gamma(\eta)s\sigma,\gamma(\rho)t\tau+i\,\gamma(\eta)t\tau)$
for all $\sigma,\tau\in S_{0}$. Since $s,t\in\text{\rm Hom}(S_{0},\mathcal{S}|_{p})$ are invertible, it follows that
[…]
This gives $\gamma(\rho)=0\in\text{\rm End}(\mathcal{S}|_{p})$ and $\gamma(\eta)=0\in\text{\rm End}(\mathcal{S}|_{p})$. Thus, $\rho=0$ and $\eta=0$. This finally implies $\omega=0$. ∎
Lemma 4.27.
Let $p\in\Sigma$ be a boundary point. Then $(N(p),T(p)\wedge\nu(p))\in\mathcal{Y}$.
Proof.
Let us fix an arbitrary point $p\in\Sigma$, and let us fix elements $\sigma,\tau\in S_{0}$. The boundary conditions for $s$ and $t$ imply that $\gamma(\nu)\circ s=s\circ\gamma_{0}(N)$ and $\gamma(\nu)\circ t=t\circ\gamma_{0}(N)$ at the point $p$. Since $\langle T,\nu\rangle_{\mathbb{R}^{n,1}}=0$, we obtain $\gamma^{\mathcal{W}}(T\wedge\nu)=\gamma^{\mathcal{W}}(T)\,\gamma^{\mathcal{W}}(\nu)$. This gives
$(s\gamma_{0}(N)\sigma,-t\gamma_{0}(N)\tau)$ $=(\gamma(\nu)s\sigma,-\gamma(\nu)t\tau)$
$=i\,\gamma^{\mathcal{W}}(T)\,\gamma^{\mathcal{W}}(\nu)\,(s\sigma,t\tau)$
$=i\,\gamma^{\mathcal{W}}(T\wedge\nu)\,(s\sigma,t\tau)$
at the point $p$. For abbreviation, let $\xi=N(p)\in\mathbb{R}^{n}$, $a=T(p)\in\mathbb{R}^{n,1}$, and $b=\nu(p)\in\mathbb{R}^{n,1}$. Then
$(s\gamma_{0}(\xi)\sigma,-t\gamma_{0}(\xi)\tau)=i\,\gamma^{\mathcal{W}}(a\wedge b)\,(s\sigma,t\tau)$
at the point $p$. Note that $(s\gamma_{0}(\xi)\sigma,-t\gamma_{0}(\xi)\tau)$ is a $\nabla^{\mathcal{W}}$-parallel section of the vector bundle $\mathcal{W}$. Moreover, it follows from Lemma 4.22 that $\gamma^{\mathcal{W}}(a\wedge b)\,(s\sigma,t\tau)$ is a $\nabla^{\mathcal{W}}$-parallel section of the vector bundle $\mathcal{W}$. Putting these facts together, we conclude that
$(s\gamma_{0}(\xi)\sigma,-t\gamma_{0}(\xi)\tau)=i\,\gamma^{\mathcal{W}}(a\wedge b)\,(s\sigma,t\tau)$
at each point in $M$. Since $\sigma,\tau\in S_{0}$ are arbitrary, it follows that $(\xi,a\wedge b)\in\mathcal{Y}$. ∎
@@
We have
$\mathcal{Y}=\{\mathbb{R}\cdot(N(p),T(p)\wedge\nu(p))\colon p\in\Sigma\}.$
The projection $\mathcal{Y}\to\mathbb{R}^{n},\,(\xi,\omega)\mapsto\xi$ is bijective. In particular, $\mathcal{Y}$ has dimension $n$.
@@
We consider the set
$\mathcal{Y}_{0}=\{\mathbb{R}\cdot(N(p),T(p)\wedge\nu(p))\colon p\in\Sigma\}\subset\mathbb{R}^{n}\times\Lambda^{2}\mathbb{R}^{n,1}.$
It follows from Lemma 4.27 that $\mathcal{Y}_{0}\subset\mathcal{Y}$. By Lemma 4.26, the projection $\mathcal{Y}\to\mathbb{R}^{n},\,(\xi,\omega)\mapsto\xi$ is injective. On the other hand, since the map $N\colon\Sigma\to S^{n-1}$ is surjective, it follows that the projection $\mathcal{Y}_{0}\to\mathbb{R}^{n},\,(\xi,\omega)\mapsto\xi$ is surjective. Putting these facts together, we conclude that $\mathcal{Y}_{0}=\mathcal{Y}$ and the projection $\mathcal{Y}\to\mathbb{R}^{n},\,(\xi,\omega)\mapsto\xi$ is bijective. ∎
@@
We have $\mathcal{Z}=\{\mathbb{R}\cdot(T(p)\wedge\nu(p))\colon p\in\Sigma\}$. Moreover, $\mathcal{Z}$ has dimension $n$.
@@
The first statement follows immediately from Proposition 4.28. The second statement follows from the fact that $\mathcal{Y}$ has dimension $n$ and the projection $\mathcal{Y}\to\Lambda^{2}\mathbb{R}^{n,1},\,(\xi,\omega)\mapsto\omega$ is injective (see Lemma 4.25). ∎
@@
It follows from Corollary 4.29 that the subspace $\mathcal{Z}\subset\Lambda^{2}\mathbb{R}^{n,1}$ is decomposable in the sense of [Ghorpade-Patil-Pillai]. (See [Ghorpade-Patil-Pillai]*p. 55 for the definition.) Applying [Ghorpade-Patil-Pillai]*Theorem 6 with $l=2$ and $r=n$, we conclude that the subspace $\mathcal{Z}\subset\Lambda^{2}\mathbb{R}^{n,1}$ is close in the sense of [Ghorpade-Patil-Pillai]. (See [Ghorpade-Patil-Pillai]*p. 57 for the definition.) Thus, there are two possibilities.
Case 1: The subspace $\mathcal{Z}\subset\Lambda^{2}\mathbb{R}^{n,1}$ is close of Type I (cf. [Ghorpade-Patil-Pillai], p. 57). In this case, there exists a basis $\{b_{0},b_{1},\ldots,b_{n}\}$ of $\mathbb{R}^{n,1}$ with the property that
$\mathcal{Z}=\text{\rm span}\{b_{0}\wedge b_{1},b_{0}\wedge b_{2},\ldots,b_{0}\wedge b_{n}\}.$
We claim that $b_{0}$ is timelike. Suppose that $b_{0}$ is spacelike or null. We can find a spacelike vector $\tilde{b}\in\mathbb{R}^{n,1}$ such that $\langle b_{0},\tilde{b}\rangle_{\mathbb{R}^{n,1}}=0$. This implies that $b_{0}$ and $\tilde{b}$ are linearly independent. Moreover, $b_{0}\wedge\tilde{b}\in\mathcal{Z}$. By Corollary 4.29, we can find a point $p\in\Sigma$ with the property that $b_{0}\wedge\tilde{b}$ is a non-zero scalar multiple of $T(p)\wedge\nu(p)$. In particular, $\text{\rm span}\{b_{0},\tilde{b}\}=\text{\rm span}\{T(p),\nu(p)\}$. This contradicts the fact that $\text{\rm span}\{b_{0},\tilde{b}\}$ does not contain a timelike vector.
Case 2: The subspace $\mathcal{Z}\subset\Lambda^{2}\mathbb{R}^{n,1}$ is close of Type II, but not close of Type I (cf. [Ghorpade-Patil-Pillai]*p. 57). In this case, $n=3$. Moreover, there exists a collection of three linearly independent vectors $b_{0},b_{1},b_{2}\in\mathbb{R}^{3,1}$ with the property that $\mathcal{Z}=\text{\rm span}\{b_{0}\wedge b_{1},b_{1}\wedge b_{2},b_{2}\wedge b_{0}\}$. We can find two linearly independent vectors $\tilde{b}_{1},\tilde{b}_{2}\in\text{\rm span}\{b_{0},b_{1},b_{2}\}$ such that $\text{\rm span}\{\tilde{b}_{1},\tilde{b}_{2}\}$ is spacelike. Then $\tilde{b}_{1}\wedge\tilde{b}_{2}\in\mathcal{Z}$. By Corollary 4.29, we can find a point $p\in\Sigma$ with the property that $\tilde{b}_{1}\wedge\tilde{b}_{2}$ is a non-zero scalar multiple of $T(p)\wedge\nu(p)$. In particular,
$\text{\rm span}\{\tilde{b}_{1},\tilde{b}_{2}\}=\text{\rm span}\{T(p),\nu(p)\}.$
This contradicts the fact that $\text{\rm span}\{\tilde{b}_{1},\tilde{b}_{2}\}$ does not contain a timelike vector. This case cannot occur. ∎
@@
For each point $p\in\Sigma$, Lemma 4.27 implies that
$T(p)\wedge\nu(p)\in\mathcal{Z}=\text{\rm span}\{b_{0}\wedge b_{1},b_{0}\wedge b_{2},\ldots,b_{0}\wedge b_{n}\};$
hence $b_{0}\in\text{\rm span}\{T(p),\nu(p)\}$. Consequently, we have $\langle b_{0},X\rangle_{\mathbb{R}^{n,1}}=0$ for every tangent vector $X\in T\Sigma$. In particular, each connected component of $\Sigma$ is contained in a hyperplane in $\mathbb{R}^{n,1}$ which is orthogonal to $b_{0}$.
For each point $p\in\Sigma$, there is a unique spacelike unit vector $B(p)\in\mathbb{R}^{n,1}$ such that $\langle b_{0},B(p)\rangle_{\mathbb{R}^{n,1}}=0$ and $T(p)\wedge\nu(p)=b_{0}\wedge B(p)$. This defines a smooth map $B$ from $\Sigma$ to the (n1)(n-1)-dimensional unit sphere
$S^{n-1}=\{b\in\mathbb{R}^{n,1}\colon\langle b_{0},b\rangle_{\mathbb{R}^{n,1}}=0,\,\langle b,b\rangle_{\mathbb{R}^{n,1}}=1\}.$
The map $B$ can be interpreted as the unit normal to $\Sigma$ within a hyperplane orthogonal to $b_{0}$.
In view of Lemma 4.26, we can find a linear transformation $A\colon\mathbb{R}^{n}\to\mathcal{Z}$ such that
$\mathcal{Y}\subset\{(\xi,A\xi)\colon\xi\in\mathbb{R}^{n}\}.$
Using Lemma 4.27, it follows that the map $\Sigma\to\mathcal{Z},\,p\mapsto T(p)\wedge\nu(p)$ is the composition of the map $N\colon\Sigma\to S^{n-1}$ with the linear transformation $A\colon\mathbb{R}^{n}\to\mathcal{Z}$. This implies that the map $B\colon\Sigma\to S^{n-1}$ is the composition of the map $N\colon\Sigma\to S^{n-1}$ with a linear isometry of $\mathbb{R}^{n}$. In particular, $\|dB\|_{\text{\rm tr}}=\|dN\|_{\text{\rm tr}}$ at each point on $\Sigma$.
@@
Suppose that $p$ is a point on the boundary $\Sigma$ with $\langle b_{0},\nu\rangle_{\mathbb{R}^{n,1}}\neq 0$. Then $h=0$, $dN=0$, and $dB=0$ at the point $p$. Moreover, the second fundamental form of $\Sigma$ in $\mathbb{R}^{n,1}$ vanishes at the point $p$.
@@
Let $\{e_{1},\ldots,e_{n-1}\}$ denote an orthonormal basis of $T_{p}\Sigma$. Differentiating the identity $T\wedge\nu=b_{0}\wedge B$ gives